OpenStudy (anonymous):

Ill give medals!!! please help :) ASAP How many solutions does the equation 6s − 3s − 9 = −2 + 3 have? Only one None Two Infinitely many

OpenStudy (welshfella):

lets try to solve it 6s -.3s - 9 = -2 + 3 3s - 9 = -1 + 3 can you continue?

OpenStudy (welshfella):

* thats -2 + 3 in the last line

OpenStudy (anonymous):

I don't really know how to solve it..

OpenStudy (anonymous):

3s-9=1...

OpenStudy (welshfella):

ok sofirst you need to get 3s on its own on left side add 9 to both sides 3x - 9 + 9 = -2 + 3 + 9 the 9's on the left disappear; 3x = -2+3 + 9 3s = 1 = 9 3s = 10

OpenStudy (welshfella):

those x should be s and lasy but one line is 3s = 1 + 9

OpenStudy (welshfella):

so if 3s = 10 what will s be ?

OpenStudy (anonymous):

soo isnt that only 1 solution?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

oh thanks do you think you can help me with another?

OpenStudy (welshfella):

sorry i've gotta go right now

OpenStudy (anonymous):

oh thank you for your help:)